Write an equation that could be used to find two consecutive even integers whose product is 24. Let x represent the first integer. Solve the equation and give the two integers.
Answer by mananth(16946)

x(x+2) = 24
x^2+2x-24=0

x^2+6x-4x-24
x(x+6)-4(x+6)=0
(x-4)(x+6)=0
x=4 OR -6